OpenStudy (anonymous):

Ok so on this what do I do (Please don't give me the answer I just want to know what I need to do on it) Medical X-rays, with a wavelength of about 10^-10m, can penetrate completely through your skin. Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1000 times the wavelength of an X-ray. find the wavelength of ultraviolet rays. Show your work.

terenzreignz (terenzreignz):

It seems too simple, but, you just do what the problem says, and, it seems it *explicitly* says '1000 times the...' which would normally indicate that you multiply by a factor of 1000.

OpenStudy (anonymous):

So I multiply 10^-10 by 1000?

terenzreignz (terenzreignz):

I would think so.

OpenStudy (anonymous):

So its 100,000,000,000,000?

terenzreignz (terenzreignz):

Okay, so, what do you think \[\Large 10^{-10}\]means?

OpenStudy (anonymous):

I did 10^-10 * 1000 and got 100,000,000,000,000

terenzreignz (terenzreignz):

How? Why? lol

OpenStudy (anonymous):

Well I asked if I multiply 10^-10 * 1000 and you said you think so :P

terenzreignz (terenzreignz):

Yes, that's true, but how exactly did you arrive at your answer? \[\Large \left.\begin{matrix}10^{-10}\cdot 1000&=&100 \ 000 \ 000 \ 000 \ 000\\&\color{red}?&\end{matrix}\right. \]

OpenStudy (anonymous):

First I multiplied out 10^-10 and got 100,000,000,000 then I multiplied it by 1000 and got 100,000,000,000,000

terenzreignz (terenzreignz):

Hang on, why is \[\Large 10^{-10}\]equal to 100 000 000 000

OpenStudy (anonymous):

I multiplied it out

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Am I not suppose to do that...

OpenStudy (klimenkov):

What is \(a^{-1}\) ?

terenzreignz (terenzreignz):

Sorry... the problem here is that the exponent was negative... so it's not as straightforward as simply 'multiplying out'

terenzreignz (terenzreignz):

But first, @klimenkov asked you a question ^_^

OpenStudy (anonymous):



terenzreignz (terenzreignz):

Okay, now what about \[\Large a^{-n}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Now can we get back to the question sorry for the rush I just have a ton of work I have to get done tonight

terenzreignz (terenzreignz):

Okay... well, if you could do that with the 'a' What is \[\Large 10^{-10}\] ?

OpenStudy (klimenkov):

Now, what is \(10^{-10}\) ?